# Better by Comparison: Retrieval-Augmented Contrastive Reasoning for Automatic Prompt Optimization

### v1

@@ IV-C2 Optimal K validation
To identify the optimal evidence of retrieved pool size in our RAG framework, we evaluate CRPO with $k\in\{5,10,15,20\}$. As shown in Figure 3, performance improves when moving from a small evidence set to a moderate pool, but declines as the pool grows further. Fixing $k=10$ yields the best trade-off between evidence diversity and inference cost, producing the most stable gains across language models and CRPO variants. This suggests the same phenomenon occurs regardless of model type, and in a limited context, overstretching the evidence tends to degrade the performance. Accordingly, we set the default Top-$K$ to $k=10$.
V Conclusion & Limitations
In this paper, we introduced Contrastive Reasoning Prompt Optimization (CRPO), a retrieval-augmented framework that improves prompt quality through tiered and multi-metric contrastive reasoning. By explicitly reflecting on high-, medium-, and low-quality prompts, as well as integrating strengths across multiple evaluation dimensions, CRPO enables LLMs to generate more robust and human-aligned outputs without fine-tuning. Experimental results on the HelpSteer2 dataset show that CRPO consistently outperforms baselines, including direct generation, CoT prompting, and simple retrieval augmentation, with ablation studies confirming the central role of contrastive reasoning in these gains.
While CRPO effectively enhances prompt optimization through tiered and multi-metric contrastive reasoning, several limitations remain. First, CRPO’s performance is tied to the HelpSteer2 dataset; while richly annotated, it may not generalize to all domains or user interaction styles encountered in practice. Second, CRPO operates in a single-turn setting, leaving its extension to multi-turn or task-specific dialogue scenarios unexplored. Third, our method leverages general-purpose retrieval (BM25), and future work could examine whether neural or hybrid retrievers provide further gains. Finally, CRPO’s evaluation depends on a reward model; thus, complementary human qualitative analysis would provide deeper insights into interpretability and practical usefulness.

### v2

@@ IV-C2 Optimal K validation
To identify the optimal evidence of retrieved pool size in our RAG framework, we evaluate CRPO with $k\in\{5,10,15,20\}$. As shown in Figure 3, performance improves when moving from a small evidence set to a moderate pool, but declines as the pool grows further. Fixing $k=10$ yields the best trade-off between evidence diversity and inference cost, producing the most stable gains across language models and CRPO variants.
V Conclusion
In this paper, we introduced Contrastive Reasoning Prompt Optimization (CRPO), a retrieval-augmented framework that improves prompt quality through tiered and multi-metric contrastive reasoning. Unlike prior methods that optimize prompts based only on prompt text and scores, CRPO explicitly reflects on high-, medium-, and low-quality prompt–response pairs, leveraging both the prompts and their generated responses to guide optimization. By integrating complementary strengths across multiple evaluation dimensions, CRPO enables LLMs to produce more robust, interpretable, and human-aligned outputs without fine-tuning. Experimental results on the HelpSteer2 dataset show that CRPO consistently outperforms baselines, including direct generation, CoT prompting, and simple retrieval augmentation, with ablation studies confirming the central role of contrastive reasoning over prompt–response pairs in achieving these gains.